OpenStudy (anonymous):

Help!!!!!: Ben opened a credit card at a department store with an APR of 27.99% compounded monthly. What is the APY on this credit card? 2.33% 31.88% 39.27% 55.98%

OpenStudy (anonymous):

\[APY = (1 + \frac{ APR }{ m })^m -1\]

OpenStudy (anonymous):

Huh..? can you break it down for me im havin a lot of trouble

OpenStudy (anonymous):

APY = (1+(APR/m))^m - 1

OpenStudy (anonymous):

m = months

OpenStudy (anonymous):

can u plug it in and explain as u go pls

OpenStudy (anonymous):

come on eulie you can do it

OpenStudy (anonymous):

i really cant im so confused and suck at math

OpenStudy (anonymous):

APY = (1+(27.99%/12))^12 - 1 APY = (1+(0.2799/12))^12 - 1 APY = (1+0.023325)^12 - 1 can you finish?

OpenStudy (anonymous):

no. i hav no idea what any of tha means

OpenStudy (anonymous):

plug in to the calculator

OpenStudy (anonymous):

mk hold on while i do tht

OpenStudy (anonymous):

oh! its 2.33 percent rite?

OpenStudy (anonymous):

no 1.023325^12 = ?

OpenStudy (anonymous):

1.318751631

OpenStudy (anonymous):

1.318751631-1= 0.31875 = 31.88%